a. What is the theoretical probability of rolling one die and getting a 5 ?
b. What is the theoretical probability of rolling one die and getting a 3 or a 4?
c. What is the theoretical probability of rolling one die and not getting a 6?
d. What is the theoretical probability of rolling one die and getting a 7 ?
Answer #1

There are 6 possible outcomes for a die: {1, 2, 3, 4, 5, 6}

Hence,

a) P(5) = 1/6

b) P(3 or 4) = (1 + 1)/6 = 2/6 = 1/3

c) P(Not getting 6) = 1 - P(6) = 1 - 1/6 = 5/6

d) P(7) = 0